OpenStudy (anonymous):

Factor the trinomial: 3x2 - 7x - 6

jimthompson5910 (jim_thompson5910):

Multiply the first and last coefficients: 3*(-6) = -18 Now find two numbers that both multiply to -18 and add to -7

jimthompson5910 (jim_thompson5910):

Which two numbers do this?

OpenStudy (anonymous):

1 & 6 ?

jimthompson5910 (jim_thompson5910):

no, 1 and 6 multiply to 6 and add to 7 The two numbers are really -9 and 2 This is because -9*2 = -18 and -9+2 = -7

jimthompson5910 (jim_thompson5910):

Now that we know that the two numbers are -9 and 2, we can break up -7x into -9x+2x So 3x^2 - 7x - 6 becomes 3x^2 - 9x + 2x - 6 Now factor by grouping 3x^2 - 9x + 2x - 6 (3x^2 - 9x) + (2x - 6) 3x(x-3) + 2(x-3) (3x+2)(x-3) This tells us that 3x^2 - 7x - 6 completely factors to (3x+2)(x-3)

OpenStudy (anonymous):

Thank you so much for your help, & explaining it out to me in a way I understand better.

jimthompson5910 (jim_thompson5910):

you're welcome